OpenStudy (anonymous):

y/x/x/3xy simply the complex fraction assume no division by 0.

OpenStudy (anonymous):

\[\frac{y \times 3xy}{x \times x} = \frac{3y^2}{x}\]

OpenStudy (anonymous):

thanks!